SPECTROSCOPIC CHARACTERIZATION OF AN ALKYL-SUBSTITUTED CRIEGEE INTERMEDIATE CH_{3}CHOO AND ITS OH RADICAL PRODUCTS

JOSEPH M. BEAMES, FANG LIU, \underline{LU LU} and MARSHA I. LESTER, Department of Chemsitry, University of Pennsylvania, Philadelphia, PA 19104-6323.

In the atmosphere, cycloaddition of ozone to the double bond of alkenes produces energized Criegee intermediates, which undergo subsequent decay processes to yield OH radicals. In this laboratory, a simple alkyl-substituted Criegee intermediate CH_{3}CHOO is produced by 248 nm photolysis of CH_{3}CHI_{2} and subsequent reaction of CH_{3}CHI with O_{2} in a quartz capillary tube reactor, following the same approach utilized for CH_{2}OO.^{a} The CH_{3}CHOO intermediate (m/z=60) and other products are detected following supersonic expansion using 118 nm VUV ionization in a time-of-flight mass spectrometer. The OH radical products from decomposition of the CH_{3}CHOO intermediate are also directly detected at m/z=17 using a new UV+VUV ionization scheme, combining UV excitation on the OH A ^{2}Σ^{+}-X ^{2}Π (1,0) transition with fixed-frequency VUV at 118 nm,^{b} or alternatively by UV laser-induced fluorescence on the OH A-X transition; OH products are also observed from CH_{2}OO. The CH_{3}CHOO intermediate is characterized by a strong B ^{1}A′-X ^{1}A′ electronic transition, in which UV excitation near the peak of a broad absorption profile centered at 320 nm results in significant depletion of the CH_{3}CHOO photoionization signal. The mechanism proposed for OH generation from energized CH_{3}CHOO and many larger Criegee intermediates is a 1,4 H-atom shift to form vinylhydroperoxide species that decay to produce OH. This reaction scheme provides a non-photolytic source of OH radicals in the atmosphere during night and winter times.

aJ. M. Beames, F. Liu, L. Lu, and M. I. Lester, J. Am. Chem. Soc. 134, 20045 (2012).

bJ. M. Beames, F. Liu, M. I. Lester and C. Murray, J. Chem. Phys. 134, 241102 (2011).